# From Neurons to Computation: Biological Reservoir Computing for Pattern Recognition

### v1

@@ Abstract
In this paper, we introduce a novel paradigm for reservoir computing (RC) that leverages a pool of cultured biological neurons as the reservoir substrate, creating a biological reservoir computing (BRC). This system operates similarly to an echo state network (ESN), with the key distinction that the neural activity is generated by a network of cultured neurons, rather than being modeled by traditional artificial computational units. The neuronal activity is recorded using a multi-electrode array (MEA), which enables high-throughput recording of neural signals. In our approach, inputs are introduced into the network through a subset of the MEA electrodes, while the remaining electrodes capture the resulting neural activity. This generates a nonlinear mapping of the input data to a high-dimensional biological feature space, where distinguishing between data becomes more efficient and straightforward, allowing a simple linear classifier to perform pattern recognition tasks effectively. To evaluate the performance of our proposed system, we present an experimental study that includes various input patterns, such as positional codes, bars with different orientations, and a digit recognition task. The results demonstrate the feasibility of using biological neural networks to perform tasks traditionally handled by artificial neural networks, paving the way for further exploration of biologically-inspired computing systems, with potential applications in neuromorphic engineering and bio-hybrid computing.
@@
Reservoir computing (RC) [39] is a powerful machine learning paradigm that utilizes a highly nonlinear dynamical system with numerous degrees of freedom to map an input space to a high-dimensional latent space. In this latent space, data samples are often more likely to be linearly separable, making subsequent classification or regression tasks more efficient. This property of RC, where complex patterns are captured through a nonlinear mapping, has led to its successful application in various domains, such as time-series prediction [3] or speech recognition [55]. One of the most prominent approaches within RC is the liquid state machine (LSM) [40], which takes advantage of a large pool of nonlinear neural units to create the desired high-dimensional feature mapping. A popular variant of this approach is the echo state network (ESN) [24, 16]. The ESN is particularly notable for its biological inspiration, as it uses biologically realistic spiking neuron models [17] that mimic the way neurons communicate and process information in the brain.
In this paper, we introduce a novel extension to the RC paradigm: the use of a cultured network of real biological neurons as the reservoir substrate. This approach, which we refer to as the biological reservoir computing (BRC), employs a biological dynamical system to map input samples into a high-dimensional space of biological features. Unlike traditional RC systems, where the reservoir is typically composed of artificial neurons or units, our biological system leverages real neurons grown in vitro, providing a natural bridge between computational and biological systems. This paradigm has the potential to improve energy efficiency – a significant limitation of current deep learning technologies [2, 25] – while also advancing our understanding of biological neuron behavior. Indeed, neuroscientific observations suggest that transient dynamics in the brain carry a significant portion of the input information, indicating that some form of RC may play a role in high-level biological cognition [13]. In more detail, we derive our neurons from stem cells using well-established neurobiological techniques [7, 8]. These cultured networks are then interfaced with a high-density multi-electrode array (MEA) device [4], which contains 4096 electrodes for both electrical stimulation and neural response recording (see also Fig. 1 for a visual overview of our approach).
@@
We introduce a novel reservoir computing paradigm that leverages biologically cultured networks as substrates for generating high-dimensional feature representations, which we term biological reservoir computing (BRC). Our approach interfaces with these cultured networks using a high-density multi-electrode array (MEA) device, enabling both electrical stimulation and the recording of neural responses.
@@ 2 Related Work
Reservoir computing (RC) [39, 49] represents one of the bio-inspired paradigms that has gained considerable attention for its potential to model complex neural dynamics. Specifically, echo state networks (ESNs) [16, 24] are a prominent subclass of liquid state machines (LSMs) [40, 56] that leverage spiking neuron models [17, 1] to implement highly nonlinear neural dynamics. In an ESN, a large pool of randomly connected spiking neurons receives an input that spreads through the network in a chaotic manner, leading to a high-dimensional, nonlinear feature space that facilitates easier separation of input patterns for downstream tasks. Although the network connectivity is random, specific algorithms are employed to ensure that the dynamics of the network are appropriately tuned to achieve effective computation [16, 48]. Several extensions of ESNs have been proposed to enhance their biological plausibility. One such extension [54] incorporates biologically plausible synaptic adaptation mechanisms based on spike-time dependent plasticity (STDP) [51, 17], which mimics how synaptic strengths are adjusted based on the timing of pre- and post-synaptic spikes. Another model, the SORN (self-organizing recurrent network) [38], integrates STDP with homeostatic plasticity [6] to maintain network stability and support learning in recurrent networks. A gating mechanism proposed in [48] further enhances the ESN by improving the recall of long-term dependencies, which are often lost in highly nonlinear recurrent systems over extended temporal ranges. A related technique, the extreme learning machine (ELM) [21, 20], utilizes a large pool of feedforward nonlinear neurons with randomly initialized weights to map inputs into a high-dimensional feature space, achieving improved linear separability despite lacking recurrent connections. While ELMs differ from RC approaches in their architecture, they share the fundamental characteristic of using random pools of nonlinear neurons to achieve a highly nonlinear mapping of input data. The RC methods described above have seen success in applications such as speech processing [56] and continual learning [11], among others.
@@
In contrast to previous work, this contribution extends the biological plausibility of RC approaches by leveraging a real biological neural network as the computational reservoir. Specifically, we introduce the concept of biological reservoir computing (BRC), where cultured biological neurons serve as the reservoir substrate. While interfacing with biological neurons through MEA devices has been explored in prior research [50, 46, 14, 23, 18, 44, 27], our study represents the first investigation into using cultured networks of biological neurons as a potential RC substrate. This novel approach offers a promising avenue for advancing the field by combining biological computation with machine learning techniques, with potential benefits including energy efficiency and a deeper understanding of biological neural dynamics.
@@ 3.2 Classifier Training Procedure
The training phase of the single-layer perceptron classifier begins by selecting a collection of input patterns to be used for network stimulation through the MEA interface. Each pattern is mapped to a predefined subset of MEA electrodes designated for stimulation. For every input, the corresponding electrodes are activated via electrical pulses repeated times to obtain a distribution of responses suitable for statistical analysis. To mitigate potential carry-over effects and ensure reliable neuronal responses, a interval is enforced between successive stimuli, allowing the cultured network to return to its resting state [46]. We set a temporal chunk – corresponding to approximately two to three synaptic transmission cycles [47] – and the number of spikes detected on each electrode within this window is counted. This results in a 4096-dimensional feature vector representing the input pattern in the latent space defined by the biological reservoir.
@@ 4 Experimental Evaluation
In this section, we present the experiments conducted to assess our BRC system and discuss the obtained results. Specifically, we evaluate the classification performance of the biological feature representations using three sets of experiments, each based on a different input pattern: (i) pointwise stimuli, (ii) oriented bars, and (iii) digits. Figure 4 illustrates the various input patterns on the MEA device. Table 4, on the other hand, summarizes the classification accuracy results. To enable a direct comparison between our biological reservoir and an artificial one, we implemented an ESN with 4,096 recurrently connected units and a connection sparsity of 10%, meaning only 10% of possible connections between units were active. To mirror our experimental setup, the ESN was initialized to a resting state with all units set to zero and driven by the same external stimulus patterns used in the biological experiments. To emulate the variability of spontaneous activity observed in the biological reservoir, we introduced input noise modeled from empirical data: we analyzed randomly selected windows of spontaneous neural activity and computed the average spike count. This estimate was used to generate realistic noise, which was superimposed on the input before being fed into the ESN. After stimulus presentation, the ESN dynamics evolved for one additional time step, and the resulting state was fed to a single-layer perceptron trained to classify the input patterns. While the ESN serves as a useful benchmark, it should be considered an upper bound, as we expect it to perform better by design. Nevertheless, our BRC system retains the distinctive advantages of a biological substrate.
@@ 4.3 Digit Recognition Experiments
Results are presented in Tab. 4. This third experimental scenario ranks between the two previous ones, with the linear classifier achieving an accuracy of 95% 6%. Although the digit recognition task is, in principle, more challenging than the oriented bar experiments, we hypothesize that the higher performance is due to the greater number of stimulated electrodes. In the oriented bar scenario, fewer electrodes are activated, likely resulting in more localized and weaker network responses. The artificial ESN competitor performs slightly better, reaching an accuracy of 97%, as expected. Nevertheless, our BRC system delivers competitive results while retaining the inherent advantages of a biological reservoir.

### v2

@@ Abstract
In this paper, we introduce a paradigm for reservoir computing (RC) that leverages a pool of cultured biological neurons as the reservoir substrate, creating a biological reservoir computing (BRC). This system operates similarly to an echo state network (ESN), with the key distinction that the neural activity is generated by a network of cultured neurons, rather than being modeled by traditional artificial computational units. The neuronal activity is recorded using a multi-electrode array (MEA), which enables high-throughput recording of neural signals. In our approach, inputs are introduced into the network through a subset of the MEA electrodes, while the remaining electrodes capture the resulting neural activity. This generates a nonlinear mapping of the input data to a high-dimensional biological feature space, where distinguishing between data becomes more efficient and straightforward, allowing a simple linear classifier to perform pattern recognition tasks effectively. To evaluate the performance of our proposed system, we present an experimental study that includes various input patterns, such as positional codes, bars with different orientations, and a digit recognition task. The results demonstrate the feasibility of using biological neural networks to perform tasks traditionally handled by artificial neural networks, paving the way for further exploration of biologically-inspired computing systems, with potential applications in neuromorphic engineering and bio-hybrid computing.
@@
Reservoir Computing (RC) [40] is a powerful machine learning paradigm that leverages a highly nonlinear dynamical system with numerous degrees of freedom to project input data into a high-dimensional latent space. In this space, data samples are often more easily linearly separable, which facilitates efficient downstream tasks such as classification or regression. Several models have been proposed to implement the reservoir, among which the Liquid State Machine (LSM) [41] and the Echo State Network (ESN) [25, 17] are the most prominent. The LSM employs a pool of nonlinear spiking neural units to generate a dynamic representation of input signals, while the ESN uses continuous-valued neurons and does not aim to replicate biologically realistic spiking behavior. These models have enabled successful applications in various domains, including time-series prediction [3] and speech recognition [57].
In this paper, we introduce an extension to the RC paradigm: the use of a cultured network of real biological neurons as the reservoir substrate. This approach, which we refer to as the biological reservoir computing (BRC), employs a biological dynamical system to map input samples into a high-dimensional space of biological features. Unlike traditional RC systems, where the reservoir is typically composed of artificial neurons or units, our biological system leverages real neurons grown in vitro, providing a natural bridge between computational and biological systems. This paradigm has the potential to improve energy efficiency – a significant limitation of current deep learning technologies [2, 26] – while also advancing our understanding of biological neuron behavior. Indeed, neuroscientific observations suggest that transient dynamics in the brain carry a significant portion of the input information, indicating that some form of RC may play a role in high-level biological cognition [14]. In more detail, we derive our neurons from stem cells using well-established neurobiological techniques [7, 8]. These cultured networks are then interfaced with a high-density multi-electrode array (MEA) device [4], which contains 4096 electrodes for both electrical stimulation and neural response recording (see also Fig. 1 for a visual overview of our approach).
@@
We introduce a reservoir computing paradigm that leverages biologically cultured networks as substrates for generating high-dimensional feature representations, which we term biological reservoir computing (BRC). Our approach interfaces with these cultured networks using a high-density multi-electrode array (MEA) device, enabling both electrical stimulation and the recording of neural responses.
@@ 2 Related Work
Reservoir Computing (RC) [40, 50] represents one of the bio-inspired paradigms that has gained considerable attention for its potential to model complex neural dynamics. Specifically, Echo State Networks (ESNs) [17, 25] and Liquid State Machines (LSMs) [41, 58] are two prominent models within the RC framework. While LSMs leverage biologically inspired spiking neuron models [18, 1] to implement highly nonlinear neural dynamics, ESNs rely on continuous-valued neurons and random recurrent connectivity to generate rich internal dynamics. In both cases, the reservoir transforms input signals into a high-dimensional feature space, facilitating pattern separation for downstream tasks. Although the connectivity is typically random, specific algorithms are employed to ensure that the network dynamics remain stable and computationally effective [17, 49]. Several extensions and related models have been proposed to enhance the biological plausibility of reservoir computing systems. For instance, some variants of ESNs [56] incorporate biologically inspired mechanisms such as spike-timing-dependent plasticity (STDP) [52, 18], which adjusts synaptic strengths based on the timing of pre- and post-synaptic activity. Another biologically motivated model, the Self-Organizing Recurrent Network (SORN) [39], combines STDP with homeostatic plasticity [6] to maintain network stability and support unsupervised learning in recurrent architectures. Additionally, a gating mechanism proposed in [49] enhances ESN performance by improving the recall of long-term dependencies, which are often degraded in highly nonlinear recurrent systems. A related technique, the Extreme Learning Machine (ELM) [22, 21], employs a large pool of feedforward nonlinear neurons with randomly initialized weights to project inputs into a high-dimensional feature space. Although ELMs lack recurrent connections and differ architecturally from RC models, they share the core principle of using random nonlinear mappings to facilitate efficient learning. These biologically inspired and computationally efficient models have demonstrated success in various applications, including speech processing [58] and continual learning [11], among others.
@@
In contrast to previous works, this contribution enhances the biological plausibility of RC approaches by leveraging a real biological neural network as the computational reservoir. Specifically, we introduce the concept of biological reservoir computing (BRC), in which cultured biological neurons serve as the reservoir substrate. Prior studies have explored interfacing with biological neurons via microelectrode array (MEA) devices [51, 47, 15, 24, 19, 45, 28], but our work specifically investigates the use of cultured neuronal networks as a potential RC substrate. A few previous efforts have also employed biological neurons in RC frameworks; however, they either focused on input separation through distinct stimuli [13] or relied on alternative interfacing methods such as optogenetics and calcium imaging instead of MEA-based systems [53]. Our approach offers a promising direction for advancing the field by integrating biological computation with machine learning methodologies, potentially yielding benefits such as improved energy efficiency and deeper insights into the dynamics of biological neural systems.
@@ 3.2 Classifier Training Procedure
The training phase of the single-layer perceptron classifier begins by selecting a collection of input patterns to be used for network stimulation through the MEA interface. Each pattern is mapped to a predefined subset of MEA electrodes designated for stimulation. For every input, the corresponding electrodes are activated via electrical pulses repeated times to obtain a distribution of responses suitable for statistical analysis. To mitigate potential carry-over effects and ensure reliable neuronal responses, a interval is enforced between successive stimuli, allowing the cultured network to return to its resting state [47]. We set a temporal chunk – corresponding to approximately two synaptic transmission cycles [48] – and the number of spikes detected on each electrode within this window is counted. This results in a 4096-dimensional feature vector representing the input pattern in the latent space defined by the biological reservoir. To test its effectiveness in transforming the input and propagating information from the stimulated neurons to the others, we excluded a square region around the stimulated area.
@@ 4 Experimental Evaluation
In this section, we present the experiments conducted to assess our BRC system and discuss the obtained results. Specifically, we evaluate the classification performance of the biological feature representations using three sets of experiments, each based on a different input pattern: (i) pointwise stimuli, (ii) oriented bars, and (iii) digits. Figure 4 illustrates the various input patterns on the MEA device. Table 4, on the other hand, summarizes the classification accuracy results. To enable a direct comparison between our biological reservoir and an artificial one, we implemented an ESN with 4,096 recurrently connected rate-based units, matching the dimensionality of the feature vectors obtained from the MEA, which consists of 4,096 electrodes. The recurrent connection matrix was initialized with a sparsity of 10%, consistent with commonly used ESN configurations and intended to reflect the sparse connectivity observed in cortical circuits. To ensure stable and rich internal dynamics, we set the spectral radius to 0.9 – a common practice in ESN design that balances dynamic richness with fading memory. Although our reservoir state does not evolve over many time steps, and long-term memory stability was not required, this setting ensures consistent and meaningful dynamics. To mirror our experimental setup, the ESN was initialized to a resting state with all units set to zero and driven by the same external stimulus patterns used in the biological experiments. To emulate the variability of spontaneous activity observed in the biological reservoir, we introduced input noise modeled from empirical data: we analyzed randomly selected windows of spontaneous neural activity and computed the average spike count. This estimate was used to generate realistic noise, which was superimposed on the input before being fed into the ESN. After stimulus presentation, the ESN dynamics evolved for one additional time step, and the resulting state was fed to a single-layer perceptron trained to classify the input patterns. While the ESN serves as a useful benchmark, it should be considered an upper bound, as we expect it to perform better by design. Nevertheless, our BRC system retains the distinctive advantages of a biological substrate.
@@ 4.3 Digit Recognition Experiments
Results are presented in Tab. 4. This third experimental scenario ranks between the two previous ones, with the linear classifier achieving an accuracy of 95% 6%. Although the digit recognition task is, in principle, more challenging than the oriented bar experiments, we hypothesize that the higher performance is due to the greater number of stimulated electrodes. In the oriented bar scenario, fewer electrodes are activated, likely resulting in more localized and weaker network responses. The artificial ESN competitor performs on average slightly better, reaching an accuracy of 97%, as expected. Nevertheless, our BRC system delivers competitive results while retaining the inherent advantages of a biological reservoir.